OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd? A.1/3 B.1/2 C.1/4 D.1/8

OpenStudy (anonymous):

can kyle repick the number that jason picked? if yes, then its \[\frac{1}{2}\times \frac{1}{2}\]if kyle cant repick what jason had picked, then \[\frac{1}{2}\times \frac{4}{9}\]